OpenStudy (anonymous):

If the cos 30° = , then the sin 60° = _____. (1 points) 0, because the angles are complementary , because the angles are complementary , because the angles are complementary 1, because the angles are complementary

OpenStudy (anonymous):

after the = sign is sqaure root 3/ 2

OpenStudy (anonymous):

and for b before the comma its 1/2

OpenStudy (anonymous):

for c before the comma its sqaure root 3/2

OpenStudy (agl202):

Correct answer is C.!!!

OpenStudy (anonymous):

thank you

OpenStudy (agl202):

Anytime!!!

Miracrown (miracrown):

It is true that cos(30) = sin(60) because the angles are complementary. It is also true that cos(30) = sin(60) = sqrt(3)/2